# ADePT: Adaptive Decomposed Prompt Tuning for Parameter-Efficient Fine-tuning

### v2

@@ 4.1 EXPERIMENTAL SETUP
Tasks and Models. We conduct extensive experiments to validate our proposed ADePT. We consider four benchmarks and 4 other datasets: (1) GLUE (Wang et al., 2018) benchmark, which includes MNLI (Williams et al., 2018), QQP [^1] https://www.quora.com/q/quoradata/, QNLI (Rajpurkar et al., 2016), SST-2 (Socher et al., 2013), STS-B (Cer et al., 2017), MRPC (Dolan & Brockett, 2005), RTE (Giampiccolo et al., 2007) and CoLA (Warstadt et al., 2019); (2) SuperGLUE benchmark (Wang et al., 2019), which includes MultiRC (Khashabi et al., 2018), BoolQ (Clark et al., 2019), WiC (Pilehvar & Camacho-Collados, 2019), WSC (Levesque et al., 2012), CB (De Marneffe et al., 2019) and ReCoRD (Zhang et al., 2018); (3) MRQA 2019 Shared Task (Fisch et al., 2019), which includes Natural Questions (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), SearchQA (Dunn et al., 2017) and NewsQA (Trischler et al., 2017); (4) MBPP benchmark (Austin et al., 2021), which is a code generation task; (5) other datasets, which includes WinoGrande (Sakaguchi et al., 2021), Yelp-2 (Zhang et al., 2015), SciTail (Khot et al., 2018) and PAWS-Wiki (Zhang et al., 2019). Following (Asai et al., 2022; Wang et al., 2023; Shi & Lipani, 2024), we evaluate our proposed ADePT on all datasets for the T5-base model (220M) (Raffel et al., 2020), except for MBPP and ReCoRD. For the T5-3B model (Raffel et al., 2020), we focus on large and challenging datasets (i.e., MNLI, ReCoRD, Natural Questions, HotpotQA, SearchQA, and NewsQA) to differentiate the performance of various PEFT methods. For the decoder-only PLMs (i.e., CodeGen-350M (Nijkamp et al., 2023) and Llama3-8B (Dubey et al., 2024)), we evaluate our proposed method ADePT on MBPP benchmark.
@@
Baselines. To evaluate our proposed ADePT, we compare it with five types of fine-tuning methods: (1) full fine-tuning (FT), which optimizes all the model parameters; (2) the vanilla PT (Lester et al., 2021), where target prompt vectors are initialized with randomly sampled top vocabularies; (3) the variants of PT using additional transfer or multi-task learning, including SPoT (Vu et al., 2022), ATTEMPT (Asai et al., 2022), and MPT (Wang et al., 2023); (4) the variants of PT using input token embedding offsets, i.e., DePT (Shi & Lipani, 2024); (5) state-of-the-art PEFT methods including Adapter(Houlsby et al., 2019), AdapterDrop (Rücklé et al., 2021), BitFit (Ben Zaken et al., 2022), HyperFomer (Karimi Mahabadi et al., 2021), HyperDecoder (Ivison & Peters, 2022), P-tuning (Liu et al., 2021), LoRA (Hu et al., 2021), LST (Sung et al., 2022), and their multi-task learning variants.
@@
Implementation Details. Following Shi & Lipani (2024), we use $100$ learnable virtual tokens as the soft prompt of PT. For our proposed ADePT, we adjust the hyperparameters to maintain an equivalent number of trainable parameters as PT. For instance, in the T5-base model, the token embedding dimension $d$ is $768$, so the number of trainable parameters is $l\times d=100\times 768=76800$. Following Shi & Lipani (2024), we search the length of soft prompt from $20$, $40$, $60$, and $80$. For ADePT, if using $60$ virtual tokens for soft prompt, the $d_{r}$ is got by solving the unequal equation $60\times 768+2\times d_{r}\times 768+d_{r}+768\leq 76800$. Thus, the $d_{r}\leq 19.49$ and $d$ is set to $19$ because the $d$ is the integer. According to this calculation method, the corresponding $d_{r}$ values for soft prompt lengths of $20$, $40$, $60$ and $80$ are $39$, $29$, $19$, and $9$, respectively. For a fair comparison of the T5-base model, we directly quote performance metrics from published papers (Mahabadi et al., 2021; Karimi Mahabadi et al., 2021; Asai et al., 2022; Wang et al., 2023; Sung et al., 2022; Shi & Lipani, 2024). For T5-3B model, we consistently use $60$ virtual tokens and bottleneck size $r=19$. Due to the lack of experimental results, for a fair comparison with the T5-3B model, we reproduce the experiments of the vanilla PT and DePT. For decoder-only PLMs, following Jain et al. (2024), we use $10$ virtual tokens for PT, $7$ virtual tokens and rank $r_{s}=3$ for DePT, $7$ virtual tokens and bottleneck size $r=1$ for ADePT, and rank $16$ for LoRA. For small datasets ($<70,000$ training samples) based on T5 model, we follow the learning strategy of Shi & Lipani (2024): we search the learning rate for the soft prompt from $13e-$ 1, $14e-$ 1, $15e-$ 1, and for the feed-forward neural network from $41e-$ 4, $51e-$ 5. For large datasets ($>70,000$ training samples) based on T5 model, we use learning rate 3e-1 for the soft prompt and $41e-$ 4 for the feed-forward neural networks. For the MBPP benchmark, following Jain et al. (2024), we use learning rates of $31e-$ 3 for the prompting-style tuning method, $41e-$ 4 for LoRA.
@@
1 sourced from (Asai et al., 2022). 2 sourced from (Sung et al., 2022). 3 sourced from (Wang et al., 2023). 4 sourced from (Shi & Lipani, 2024). (m) represents additional multi-task training. (t) represents additional transfer learning.
Table 4: The experimental results on MRQA 2019 Shared Task and other datasets with the associated size of trainable parameters. All results are based on the T5-base model. We report the F1 for MRQA tasks and accuracy for other datasets as test metrics. The results are averaged over three runs and the subscripts denote standard deviation. All baseline results are quoted from (Shi & Lipani, 2024).
@@
MPT 77.6K $72.0_{0.1}$ $75.8_{0.1}$ $77.2_{0.1}$ $63.7_{0.1}$ $72.2$ $56.5_{0.9}$ $96.4_{0.0}$ $95.5_{0.1}$ $93.5_{0.1}$ $85.5$
DePT 76.8K $73.2_{0.1}$ $76.8_{0.3}$ $77.6_{0.2}$ $64.4_{0.1}$ $73.0$ $59.0_{0.2}$ $96.8_{0.1}$ $95.6_{0.2}$ ${93.7}_{0.1}$ $86.3$
ADePT (ours) 76.1K ${73.9}_{0.0}$ ${77.1}_{0.1}$ ${78.7}_{0.1}$ ${64.7}_{0.1}$ 73.673.6{73.6}73.6 ${59.1}_{0.9}$ ${96.8}_{0.0}$ ${95.9}_{0.3}$ $93.7_{0.2}$ 86.486.4{86.4}86.4
@@ 4.4 Results based on Decoder-only PLMs
We evaluate our proposed ADePT on decoder-only PLMs (i.e., CodeGen-350M model (Nijkamp et al., 2023) and Llama3-8B model (Dubey et al., 2024)) through instruction tuning (Ouyang et al., 2022). We use MBPP benchmark, which is a Python program generation task (Austin et al., 2021). Following Jain et al. (2024), we use a 50-50 split for training and test. We report average $\textit{pass}@1$ scores to evaluate the performance, as shown in Table 7. We can observe that ADePT performs best among PT-style methods, demonstrating its effectiveness. In the Llama3-8B model, the vanilla PT performs much worse than DePT and ADePT. This shows that the inability of PT to change the relative attention patterns limits its adaptation ability, whereas DePT and ADePT perform better because they can change the relative attention patterns. Also, in CodeGen-350M and Llama3-8B, the use of adaptive token embedding offsets helps ADePT perform better than DePT. Although ADePT performs slightly worse than LoRA on the MBPP benchmark, ADePT requires far fewer parameters than LoRA. More importantly, LoRA needs to merge weights, which typically limits it to only a single downstream task. In contrast, ADePT can flexibly switch between tasks and adapt to multiple downstream tasks simultaneously, which is a unique advantage of PT-style methods.
@@ Appendix A How Decomposed Prompt Tuning Affect the first multi-head self-attention layer?
$&\bm{o}_{i}^{\text{DePT}}=\text{Attention}\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}+\Delta\bm{E}]\bm{W}_{K}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}+\Delta\bm{E}]\bm{W}_{V}^{\mathcal{H}}\right),\\ &=\sum_{k=1}^{l}A_{ik}\bm{p}_{k}\bm{W}_{V}^{\mathcal{H}}+\left(1-\sum_{k=1}^{l}A_{ik}\right)\text{Softmax}\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{E}+\Delta\bm{E}\right)\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)\Delta\bm{E}\bm{W}_{V}^{\mathcal{H}}\\ &+(1-\sum_{k=1}^{l}A_{ik})\text{Softmax}\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{E}+\Delta\bm{E}\right)\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)\bm{E}\bm{W}_{V}^{\mathcal{H}},\\ &\text{with}\quad A_{ik}=\frac{\exp\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{\top}\right)}{C},\\ &C=\sum_{k=1}^{l}\exp\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{\top}\right)\\ &+\sum_{j=1}^{s}\exp\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{e}_{j}+\Delta\bm{e}_{j}\right)\bm{W}_{K}^{\mathcal{H}}\right)^{\top}\right),\end{aligned}$ (8)
@@ B.2 Few-shot Learning
Table 15: Few-shot learning results, obtained from three runs, with $k$ = {4, 16, 32} training samples on the BooQ, CB and SciTail datasets. Baseline results are directly quoted from Shi & Lipani (2024).
@@ B.3 Unsuccessful Attempts
We attempted to perform instruction tuning on a larger scale of the GSM8K dataset (Cobbe et al., 2021) using Llama3 8B, with 7,473 examples in the training set and 1,319 examples in the test set. The goal of this instruction tuning was to fine-tune the model to better solve mathematical reasoning problems in the GSM8K dataset. However, unfortunately, we were not successful in our training attempts, including both ADePT and DePT. This might be because the simple low-rank matrix multiplication and shallow neural networks are unable to fit the input token embedding offsets required by large datasets when working with large models that have a very large vocabulary. Addressing this issue will be a direction for our future research.
@@ Appendix C Additional Implementation Details
We implement our experiments by using Pytorch [^2] https://pytorch.org/, Huggingface Transformers [^3] https://github.com/huggingface/transformers, and Huggingface PEFT  [^4] https://github.com/huggingface/peft. We evaluate our proposed ADePT in four PLMs, i.e., T5-base model [^5] https://huggingface.co/google-t5/t5-base, T5-3B model [^6] https://huggingface.co/google-t5/t5-3b, CodeGen-350M [^7] https://huggingface.co/Salesforce/codegen-350M-mono and Llama3-8B [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B. Following Asai et al. (2022); Wang et al. (2023); Shi & Lipani (2024), we train the T5 model using the original checkpoint rather than the LM-adapted 1.1 version (Lester et al., 2021). For the T5-3B model, due to the limitations of computational resources, we select the several most convincing datasets to evaluate our proposed ADePT. We think that convincing datasets should have a large training dataset and large test dataset, and be challenging for the T5-base model. We use the criteria of more than 70,000 training samples, an accuracy/F1 score of less than 90% on the T5-base model, and more than 4,000 test samples to select the datasets to test our proposed ADePT on the T5-3B model. Shi & Lipani (2024) found that training PT for additional steps typically leads to performance improvements, and we follow this setting to train our proposed ADePT. We measure the latency of ADePT by running a feed-forward neural network for each token in real time.
@@
For the small datasets ($<70,000$ training samples), following Shi & Lipani (2024), we search the learning rates for the soft prompt from {$13e-1,4e-1,5e-$ 1} and for the feed-forward neural network from {$51e-4,1e-$ 5}. We also search for the prompt length from {$20,40,60,80$} with corresponding bottleneck sizes of {$39,29,19,9$} to ensure the number of trainable parameters remains below 76.8K. For the large datasets ($>70,000$ training samples), we set the prompt length as $60$, the bottleneck size as $19$, the prompt learning rate as $13e-$ 1, and the feed-forward neural network learning rate as $41e-$ 4. For decoder-only PLMs, following Jain et al. (2024), we use $10$ virtual tokens for PT, $7$ virtual tokens and rank $r_{s}=3$ for DePT, $7$ virtual tokens and bottleneck size $r=1$ for ADePT, and rank $16$ for LoRA. For the MBPP benchmark, following Jain et al. (2024), we use learning rates of $31e-$ 3 for the prompting-style tuning method and $41e-$ 4 for LoRA.
@@
Following Shi & Lipani (2024), for the T5-base model and the small datasets, we train the model for 30,000 steps; for the T5-base model and the large datasets, we train the model for 300,000 steps. For the T5-3B model, we train the model for 30,000 steps. In each trial of the t5-base model and the T5-3B model, we evaluate the performance every 1,000 steps and select the best checkpoint based on the optimal performance on the evaluation set. For the MBPP benchmarks, following Jain et al. (2024), we train the model for 10 epochs. We train the model with a batch size of 32, except for the MBPP benchmark, where we use a batch size of 4. We typically use a maximum sequence length of 256, except for SuperGLUE-MultiRC, where the maximum sequence length is 348, and MRQA, where it is 512.
@@
For the few-shot learning, following the prior works (Asai et al., 2022; Wang et al., 2023; Shi & Lipani, 2024), we first pre-train five source tasks (i.e., MNLI, QQP, SST-2, SQUAD, and ReCoRD), and then select the best checkpoint to use as the initialization for few-shot training.

### v3

@@ 4.1 EXPERIMENTAL SETUP
Tasks and Models. We conduct extensive experiments to validate our proposed ADePT. We consider four benchmarks and 4 other datasets: (1) GLUE (wang-etal-2018-glue) benchmark, which includes MNLI (williams-etal-2018-broad), QQP [^1] https://www.quora.com/q/quoradata/, QNLI (rajpurkar-etal-2016-squad), SST-2 (socher-etal-2013-recursive), STS-B (cer-etal-2017-semeval), MRPC (dolan-brockett-2005-automatically), RTE (giampiccolo-etal-2007-third) and CoLA (warstadt2019neural_cola); (2) SuperGLUE benchmark (wang2019superglue), which includes MultiRC (khashabi-etal-2018-looking), BoolQ (clark-etal-2019-boolq), WiC (pilehvar-camacho-collados-2019-wic), WSC (levesque2012winograd_wnli), CB (de2019commitmentbank) and ReCoRD (zhang2018record); (3) MRQA 2019 Shared Task (fisch-etal-2019-mrqa), which includes Natural Questions (kwiatkowski-etal-2019-natural), HotpotQA (yang-etal-2018-hotpotqa), SearchQA (dunn2017searchqa) and NewsQA (trischler-etal-2017-newsqa); (4) MBPP benchmark (MBPP), which is a code generation task; (5) other datasets, which includes WinoGrande (10.1145/3474381), Yelp-2 (10.5555/2969239.2969312), SciTail (Khot_Sabharwal_Clark_2018) and PAWS-Wiki (zhang-etal-2019-paws). Following (asai-etal-2022-attempt; MPT; shidept), we evaluate our proposed ADePT on all datasets for the T5-base model (220M) (t5), except for MBPP and ReCoRD. For the T5-3B model (t5), we focus on large and challenging datasets (i.e., MNLI, ReCoRD, Natural Questions, HotpotQA, SearchQA, and NewsQA) to differentiate the performance of various PEFT methods. For the decoder-only PLMs (i.e., CodeGen-350M (nijkamp2023codegen) and Llama3-8B (llama3)), we evaluate our proposed method ADePT on MBPP benchmark.
@@
Baselines. To evaluate our proposed ADePT, we compare it with five types of fine-tuning methods: (1) full fine-tuning (FT), which optimizes all the model parameters; (2) the vanilla PT (lester-etal-2021-power), where target prompt vectors are initialized with randomly sampled top vocabularies; (3) the variants of PT using additional transfer or multi-task learning, including SPoT (vu-etal-2022-spot), ATTEMPT (asai-etal-2022-attempt), and MPT (MPT); (4) the variants of PT using input token embedding offsets, i.e., DePT (shidept); (5) state-of-the-art PEFT methods including Adapter(houlsby2019parameter), AdapterDrop (ruckle-etal-2021-adapterdrop), BitFit (zaken2021bitfit), HyperFomer (mahabadi2021parameter), HyperDecoder (ivison-peters-2022-hyperdecoders), P-tuning (liu2021gpt), LoRA (hu2021lora), LST (sunglst), and their multi-task learning variants.
@@
Implementation Details. Following shidept, we use $100$ learnable virtual tokens as the soft prompt of PT. For our proposed ADePT, we adjust the hyperparameters to maintain an equivalent number of trainable parameters as PT. For instance, in the T5-base model, the token embedding dimension $d$ is $768$, so the number of trainable parameters is $l\times d=100\times 768=76800$. Following shidept, we search the length of soft prompt from $20$, $40$, $60$, and $80$. For ADePT, if using $60$ virtual tokens for soft prompt, the $d_{r}$ is got by solving the unequal equation $60\times 768+2\times d_{r}\times 768+d_{r}+768\leq 76800$. Thus, the $d_{r}\leq 19.49$ and $d$ is set to $19$ because the $d$ is the integer. According to this calculation method, the corresponding $d_{r}$ values for soft prompt lengths of $20$, $40$, $60$ and $80$ are $39$, $29$, 1919, and $9$, respectively. For a fair comparison of the T5-base model, we directly quote performance metrics from published papers (karimi2021compacter; mahabadi2021parameter; asai-etal-2022-attempt; MPT; sunglst; shidept). For T5-3B model, we consistently use $60$ virtual tokens and bottleneck size $r=19$. Due to the lack of experimental results, for a fair comparison with the T5-3B model, we reproduce the experiments of the vanilla PT and DePT. For decoder-only PLMs, following promptback, we use $10$ virtual tokens for PT, $7$ virtual tokens and rank $r_{s}=3$ for DePT, $7$ virtual tokens and bottleneck size $r=1$ for ADePT, and rank $16$ for LoRA. For small datasets (<70,000<70,000 training samples) based on T5 model, we follow the learning strategy of shidept: we search the learning rate for the soft prompt from 3e13e-1, 4e14e-1, 5e15e-1, and for the feed-forward neural network from 1e41e-4, 1e51e-5. For large datasets (>70,000>70,000 training samples) based on T5 model, we use learning rate 3e-1 for the soft prompt and 1e41e-4 for the feed-forward neural networks. For the MBPP benchmark, following promptback, we use learning rates of 1e31e-3 for the prompting-style tuning method, 1e41e-4 for LoRA.
@@
1 sourced from (asai-etal-2022-attempt). 2 sourced from (sunglst). 3 sourced from (MPT). 4 sourced from (shidept). (m) represents additional multi-task training. (t) represents additional transfer learning.
Table 4: The experimental results on MRQA 2019 Shared Task and other datasets with the associated size of trainable parameters. All results are based on the T5-base model. We report the F1 for MRQA tasks and accuracy for other datasets as test metrics. The results are averaged over three runs and the subscripts denote standard deviation. All baseline results are quoted from (shidept).
@@
MPT 77.6K $72.0_{0.1}$ $5.8_{0.1} 77$$7.2_{0.1}$ $63.7_{0.1}$ $72.2$ $56.5_{0.9}$ $96.4_{0.0}$ $95.5_{0.1}$ $93.5_{0.1}$ $85.5$
DePT 76.8K $73.2_{0.1}$ $6.8_{0.3} 77$$7.6_{0.2}$ $64.4_{0.1}$ $73.0$ $59.0_{0.2}$ $96.8_{0.1}$ $95.6_{0.2}$ ${93.7}_{0.1}$ $86.3$
ADePT (ours) 76.1K ${73.9}_{0.0}$ ${77.1}_{0.1}$ ${78.7}_{0.1}$ ${64.7}_{0.1}$ 73.6{73.6} ${59.1}_{0.9}$ ${96.8}_{0.0}$ ${95.9}_{0.3}$ $93.7_{0.2}$ 86.4{86.4}
@@ 4.4 Results based on Decoder-only PLMs
We evaluate our proposed ADePT on decoder-only PLMs (i.e., CodeGen-350M model (nijkamp2023codegen) and Llama3-8B model (llama3)) through instruction tuning (ouyang2022training). We use MBPP benchmark, which is a Python program generation task (MBPP). Following promptback, we use a 50-50 split for training and test. We report average $\textit{pass}@1$ scores to evaluate the performance, as shown in Table 7. We can observe that ADePT performs best among PT-style methods, demonstrating its effectiveness. In the Llama3-8B model, the vanilla PT performs much worse than DePT and ADePT. This shows that the inability of PT to change the relative attention patterns limits its adaptation ability, whereas DePT and ADePT perform better because they can change the relative attention patterns. Also, in CodeGen-350M and Llama3-8B, the use of adaptive token embedding offsets helps ADePT perform better than DePT. Although ADePT performs slightly worse than LoRA on the MBPP benchmark, ADePT requires far fewer parameters than LoRA. More importantly, LoRA needs to merge weights, which typically limits it to only a single downstream task. In contrast, ADePT can flexibly switch between tasks and adapt to multiple downstream tasks simultaneously, which is a unique advantage of PT-style methods.
@@ Appendix A How Decomposed Prompt Tuning Affect the first multi-head self-attention layer?
$\begin{aligned} &\bm{o}_{i}^{\text{DePT}}=\text{Attention}\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}+\Delta\bm{E}]\bm{W}_{K}^{\mathcal{H}},\text{concat}[\bm{P},\bm{E}+\Delta\bm{E}]\bm{W}_{V}^{\mathcal{H}}\right),\\ &=\sum_{k=1}^{l}A_{ik}\bm{p}_{k}\bm{W}_{V}^{\mathcal{H}}+\left(1-\sum_{k=1}^{l}A_{ik}\right)\text{Softmax}\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{E}+\Delta\bm{E}\right)\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)\Delta\bm{E}\bm{W}_{V}^{\mathcal{H}}\\ &+(1-\sum_{k=1}^{l}A_{ik})\text{Softmax}\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{E}+\Delta\bm{E}\right)\bm{W}_{K}^{\mathcal{H}}\right)^{T}\right)\bm{E}\bm{W}_{V}^{\mathcal{H}},\\ &\text{with}\quad A_{ik}=\frac{\exp\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{\top}\right)}{C},\\ &C=\sum_{k=1}^{l}\exp\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\bm{p}_{k}\bm{W}_{K}^{\mathcal{H}}\right)^{\top}\right)\\ &+\sum_{j=1}^{s}\exp\left(\left(\left(\bm{e}_{i}+\Delta\bm{e}_{i}\right)\bm{W}_{Q}^{\mathcal{H}}\right)\left(\left(\bm{e}_{j}+\Delta\bm{e}_{j}\right)\bm{W}_{K}^{\mathcal{H}}\right)^{\top}\right),\end{aligned}$ (8)
@@ B.2 Few-shot Learning
Table 15: Few-shot learning results, obtained from three runs, with $k$ = {4, 16, 32} training samples on the BooQ, CB and SciTail datasets. Baseline results are directly quoted from shidept.
@@ B.3 Unsuccessful Attempts
We attempted to perform instruction tuning on a larger scale of the GSM8K dataset (gsm8k) using Llama3 8B, with 7,473 examples in the training set and 1,319 examples in the test set. The goal of this instruction tuning was to fine-tune the model to better solve mathematical reasoning problems in the GSM8K dataset. However, unfortunately, we were not successful in our training attempts, including both ADePT and DePT. This might be because the simple low-rank matrix multiplication and shallow neural networks are unable to fit the input token embedding offsets required by large datasets when working with large models that have a very large vocabulary. Addressing this issue will be a direction for our future research.
@@ Appendix C Additional Implementation Details
We implement our experiments by using Pytorch [^2] https://pytorch.org/, Huggingface Transformers [^3] https://github.com/huggingface/transformers, and Huggingface PEFT  [^4] https://github.com/huggingface/peft. We evaluate our proposed ADePT in four PLMs, i.e., T5-base model [^5] https://huggingface.co/google-t5/t5-base, T5-3B model [^6] https://huggingface.co/google-t5/t5-3b, CodeGen-350M [^7] https://huggingface.co/Salesforce/codegen-350M-mono and Llama3-8B [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B. Following asai-etal-2022-attempt; MPT; shidept, we train the T5 model using the original checkpoint rather than the LM-adapted 1.1 version (lester-etal-2021-power). For the T5-3B model, due to the limitations of computational resources, we select the several most convincing datasets to evaluate our proposed ADePT. We think that convincing datasets should have a large training dataset and large test dataset, and be challenging for the T5-base model. We use the criteria of more than 70,000 training samples, an accuracy/F1 score of less than 90% on the T5-base model, and more than 4,000 test samples to select the datasets to test our proposed ADePT on the T5-3B model. shidept found that training PT for additional steps typically leads to performance improvements, and we follow this setting to train our proposed ADePT. We measure the latency of ADePT by running a feed-forward neural network for each token in real time.
@@
For the small datasets (<70,000<70,000 training samples), following shidept, we search the learning rates for the soft prompt from {3e1,4e1,5e13e-1,4e-1,5e-1} and for the feed-forward neural network from {1e4,1e51e-4,1e-5}. We also search for the prompt length from {20,40,60,8020,40,60,80} with corresponding bottleneck sizes of {39,29,19,939,29,19,9} to ensure the number of trainable parameters remains below 76.8K. For the large datasets (>70,000>70,000 training samples), we set the prompt length as $60$, the bottleneck size as $19$, the prompt learning rate as 3e13e-1, and the feed-forward neural network learning rate as 1e41e-4. For decoder-only PLMs, following promptback, we use $10$ virtual tokens for PT, $7$ virtual tokens and rank $r_{s}=3$ for DePT, $7$ virtual tokens and bottleneck size $r=1$ for ADePT, and rank $16$ for LoRA. For the MBPP benchmark, following promptback, we use learning rates of 1e31e-3 for the prompting-style tuning method and 1e41e-4 for LoRA.
@@
Following shidept, for the T5-base model and the small datasets, we train the model for 30,000 steps; for the T5-base model and the large datasets, we train the model for 300,000 steps. For the T5-3B model, we train the model for 30,000 steps. In each trial of the t5-base model and the T5-3B model, we evaluate the performance every 1,000 steps and select the best checkpoint based on the optimal performance on the evaluation set. For the MBPP benchmarks, following promptback, we train the model for 10 epochs. We train the model with a batch size of 32, except for the MBPP benchmark, where we use a batch size of 4. We typically use a maximum sequence length of 256, except for SuperGLUE-MultiRC, where the maximum sequence length is 348, and MRQA, where it is 512.
@@
For the few-shot learning, following the prior works (asai-etal-2022-attempt; MPT; shidept), we first pre-train five source tasks (i.e., MNLI, QQP, SST-2, SQUAD, and ReCoRD), and then select the best checkpoint to use as the initialization for few-shot training.